I'm wondering how 2 or 2.00 Divided by 0.004 is equal to 500. Thanks
Answer by richwmiller(17219)

2 = 2000/1000
.004 =4/1000
2000/1000 divided by 4/1000 equals
2000/4
2000/4=500